OpenStudy (anonymous):

If f(x) = x2 + 5 and g(x) = 5x + 4, what is [f o g](6)? a. 34 b. 209 c. 1,161 d. 1,394

OpenStudy (anonymous):

First find g(6)

OpenStudy (anonymous):

g(6) = 5(6) + 4 = 34

OpenStudy (anonymous):

Now find f(34) = 34^2 + 5 = 1156 + 5 = 1161

OpenStudy (anonymous):

Thank you!